# Reasoning Models Better Express Their Confidence

### v1

@@ Abstract
Despite their strengths, large language models (LLMs) often fail to communicate their confidence accurately, making it difficult to assess when they might be wrong and limiting their reliability. In this work, we demonstrate that reasoning models—LLMs that engage in extended chain-of-thought (CoT) reasoning—exhibit superior performance not only in problem-solving but also in accurately expressing their confidence. Specifically, we benchmark six reasoning models across six datasets and find that they achieve strictly better confidence calibration than their non-reasoning counterparts in 33 out of the 36 settings. Our detailed analysis reveals that these gains in calibration stem from the slow thinking behaviors of reasoning models—such as exploring alternative approaches and backtracking—which enable them to adjust their confidence dynamically throughout their CoT, making it progressively more accurate. In particular, we find that reasoning models become increasingly better calibrated as their CoT unfolds, a trend not observed in non-reasoning models. Moreover, removing slow thinking behaviors from the CoT leads to a significant drop in calibration. Lastly, we show that these gains are not exclusive to reasoning models—non-reasoning models also benefit when guided to perform slow thinking via in-context learning. [^1] Our code is available at https://github.com/MattYoon/reasoning-models-confidence
1 Introduction
[…]
We then conduct a detailed analysis to support our claim that the calibration gains in reasoning models stem from slow thinking, which allows them to dynamically adjust and correct their confidence throughout the reasoning process. (1) We measure changes in calibration metrics throughout the CoT process and observe a steady, gradual improvement for reasoning models, with statistically significant trends ($p<0.05$). Non-reasoning models show no significant trend—and in some cases, surprisingly, even exhibit a worsening pattern. (2) In our ablation study, we find that the non-linear structure of slow thinking gandhi2025cognitive —such as the model’s ability to explore alternatives and revise its reasoning—plays a critical role in improving calibration. Conversely, prompting the model to explicitly reason about its own confidence yields only limited gains in calibration. (3) Lastly, we observe that non-reasoning models can also benefit from slow thinking when guided via in-context learning (dong2022survey, ), supporting the view that these improvements arise from the slow thinking process itself, rather than from something inherently exclusive to reasoning models.
@@ 1 Introduction
Figure 1: R1-Distill-Qwen-32B dynamically refines its confidence throughout CoT (left) as it engages in various slow thinking behaviors (right). We collect the model’s answer and confidence at each token position by appending “</think>Answer:” to terminate the reasoning process early. Note that the model’s answer is consistent and correct (“Cleisthenes”) at all points, while the confidence fluctuates. See Appendix A.4 for the full untruncated CoT.
@@ 2 Related Work
Various approaches have been proposed for confidence estimation—the process of inferring an LLM’s confidence in its predictions (geng-etal-2024-survey, ). This includes measuring token-level probabilities of answers (gupta2024language, ), training proxy probe models on internal hidden states (kadavath2022language, ), and applying supervised fine-tuning with labeled confidence data (xu2024sayself, ). Nonetheless, these methods assume access to a model’s internal states or weights, which are usually unavailable in proprietary LLMs (shrivastava2023llamas, ; xie2024survey, ). Additionally, they are often model-specific, necessitating readjustment or retraining when the underlying LLM changes (yang2024verbalized, ). An alternative approach that avoids these limitations involves sampling multiple responses for the same prompt and deriving confidence based on response consistency (manakul-etal-2023-selfcheckgpt, ). While this method is compatible with proprietary LLMs and model-agnostic, its primary limitation is the increased computational cost due to repeated inference (yang2024verbalized, ).
In this work, we focus on verbalized confidence estimation, where the LLM is prompted to directly express its confidence as a part of their output, either through a linguistic phrase (e.g., “Highly likely”) or a numerical probability (e.g., “0.95”) (geng-etal-2024-survey, ). Verbalized confidence estimation has emerged as a widely studied approach as it is compatible with proprietary LLMs, model-agnostic, and computationally efficient (tian-etal-2023-just, ; xiong2024can, ; yang2024verbalized, ). While verbalized confidence estimation offers ease of use and broad applicability, its accessibility also makes its limitations more consequential. Overconfidence is a persistent issue: LLMs tend to express high certainty even when they are incorrect, posing serious risks for real-world deployment of LLMs (xiong2024can, ; zhou2024relying, ). Our results suggest that reasoning models—and their use of slow thinking—hold strong potential to mitigate this issue.
@@
Increasing compute at inference time, or inference-time scaling, has become a focus for improving the reasoning performance of LLMs (snell2025scaling, ). One effective approach to inference-time scaling is training models to generate longer CoTs, as demonstrated by models like OpenAI’s o1 (jaech2024openai, ) and DeepSeek-R1 (deepseekr1, ), which are commonly referred to as reasoning models. The key factor that distinguishes the CoT of reasoning models from that of non-reasoning, instruction-tuned models—beyond just length—is their slow thinking (min2024imitate, ), which drives their problem-solving ability. Slow thinking mirrors human cognitive behaviors, characterized by non-linear reasoning traces such as exploring alternative approaches, verifying answers, and backtracking (gandhi2025cognitive, ). During slow thinking, reasoning models also frequently produce epistemic markers that signal uncertainty—such as “I think”, and “maybe”—which are rarely observed in non-reasoning models (zhou2024relying, ). Structurally, reasoning models typically enclose their CoT within special tokens such as <think>...</think> to isolate the reasoning phase, in contrast to non-reasoning models.
@@ 3 Reasoning models better express their confidence
In this section, we demonstrate that reasoning models express their confidence more accurately than non-reasoning models, first detailing the experimental setup (Section 3.1) and then presenting the results (Section 3.2). Full results and additional experiments are provided in Appendix A, and setup details are provided in Appendix B.
@@ 3.1 Experiment setup
We evaluate LLM calibration across two types of datasets: knowledge-focused and reasoning-intensive. We use the knowledge-focused datasets—TriviaQA and NonambigQA (a nonambiguous subset of NQ-Open) (joshi-etal-2017-triviaqa, ; min-etal-2020-ambigqa, ; kwiatkowski-etal-2019-natural, )— to provide a more controlled setting for comparing the calibration of reasoning and non-reasoning models, as both achieve similar accuracy given that CoT does little to help solve these tasks (sprague2024cot, ). Therefore, this setup ensures that any differences in calibration are not simply due to reasoning models being better at solving the task. In addition, since reasoning models are primarily intended for complex reasoning tasks, we also include reasoning-intensive datasets—MMLU-Pro and SuperGPQA (mmlupro, ; supergpqa, )— to ensure our findings generalize. These datasets are challenging multiple-choice benchmarks centered on knowledge-driven reasoning. We use two subsets of each reasoning dataset: a Math subset focused on arithmetic reasoning, and a Non-Math subset covering other types of reasoning. For our experiments, we uniformly sample 1,000 examples from each dataset and subset. To assess variability, we perform bootstrapping over multiple resampled subsets and report the standard deviation in Appendix A.2.4.
@@
We benchmark a diverse set of six reasoning models derived from four different 32B-scale LLMs against their non-reasoning counterparts. [^2] We discuss the effect of model scale in Section 5. Specifically, we evaluate: R1-Distill-Qwen, QwQ, and OR1-Preview reasoning models (deepseekr1, ; qwq32b, ; skywork-or1-2025, ) against Qwen2.5-Instruct (qwen25technicalreport, ); GLM-Z1-0414 against GLM-4-0414 (glm2024chatglm, ); EXAONE-Deep (exaonedeep, ) against EXAONE-3.5-Instruct (exaone35, ); and Qwen3 with Thinking Mode (abbreviated as Qwen3 Thinking) against Qwen3 with Non-thinking Mode (Qwen3 Non-thinking) (qwen3, ). Unlike the other pairs, which consist of separate model checkpoints, Qwen3 is a hybrid model that supports both reasoning-style CoT (Thinking Mode) and non-reasoning-style CoT (Non-Thinking Mode) via prompting. [^3] The Non-thinking mode is enabled by injecting an empty thinking block, <think></think>, at the beginning of the model’s response.
@@
In a single turn of conversation, we instruct the models to perform three steps using CoT: (1) Solution Reasoning, where the model reasons step-by-step to arrive at an answer to the given question; (2) Confidence Reasoning, where it evaluates its own confidence for that answer step-by-step; and (3) Confidence Verbalization, where it maps their confidence in one of ten bins, ranging from “Almost no chance (0–0.1)” to “Almost certain (0.9–1.0)”. Each bin includes both a linguistic descriptor (e.g., “Almost certain”) and its corresponding numerical probability (e.g., “0.9–1.0”). To ensure a fair comparison, we use the same instructions for both reasoning and non-reasoning models. The full prompt we use is included in Appendix B.2.
For reasoning models, we expect all three steps to be carried out within the thinking process <think>...</think>. However, we observe that some reasoning models—specifically R1-Distill, OR1-Preview, and GLM-Z1—rarely engage in Confidence Reasoning. [^4] In other words, their thinking process mostly only contains Solution Reasoning, although they consistently do output their Confidence Verbalization outside of the thinking termination token </think> without CoT. To address this, we force these models to include Confidence Reasoning within their thinking process by generating up to the </think> token, and replacing the token with “Okay, now let’s assess my overall thinking process so far step-by-step. I need to evaluate how likely my answer is correct.” and performing a second round of inference. We ablate the effect of this choice in Section 4.2.
As part of the instruction, models are asked to format their final response as “Answer:ANSWER Confidence:CONFIDENCE” after completing all three reasoning steps. We apply rule-based filtering using regular expressions to extract the predicted answer and confidence. Answers are matched against the ground truth using GPT-4o mini (openai_gpt4omini, ), leveraging the prompt from OpenAI’s simple-evals codebase (openai_simpleevals, ). We describe the full procedure we use to ensure that all model responses contain extractable answers and confidence predictions in Appendix B.2.
@@
Expected Calibration Error (ECE(naeini2015obtaining, ) measures the average discrepancy between accuracy and predicted confidence within each confidence bin, weighted by the number of samples in each bin. While ECE is an intuitive metric for measuring absolute calibration, it fails to capture calibration at the individual prediction level and does not account for the model’s discriminative ability, which refers to how well the model assigns higher confidence to correct predictions over incorrect ones (geng-etal-2024-survey, ). In contrast, AUROC (bradley1997use, ) captures the model’s discriminative ability by computing the probability that a randomly chosen correct prediction is assigned higher confidence than a randomly chosen incorrect one, but it does not measure absolute calibration. Brier Score (brier1950verification, ) quantifies the mean squared difference between predicted confidence and the true binary outcome, capturing both absolute calibration at the individual level and the discriminative ability.
Expanded setup for robustness check
To assess the robustness of our findings, we experiment with a wide range of alternative setups, as detailed in Appendix A.2. (1) We explore different confidence expression styles—e.g., providing only a linguistic descriptor without a probability, or outputting a numerical probability directly without binning. (2) Since some reasoning models perform two rounds of inference due to the forced Confidence Reasoning, we also evaluate a similar two-step sequential prompting setup for non-reasoning models.
(3) We evaluate more advanced prompting strategies for non-reasoning models, as suggested by prior work, such as Top-K and Multi-step prompting (tian-etal-2023-just, ; xiong2024can, ). (4) While our main experiments use greedy decoding to avoid randomness and ensure reproducibility, we also test sampling-based decoding to confirm that our results generalize across decoding strategies. Across all setups, our findings remain consistent: reasoning models exhibit better calibration than their non-reasoning counterparts.
@@ 3.2 Experiment result
We observe that reasoning models consistently outperform their non-reasoning counterparts across all calibration metrics on knowledge-focused datasets (Table 1). Notably, reasoning models exhibit superior calibration even in cases where they underperform non-reasoning models in task accuracy. For example, GLM-Z1-0414 achieves 0.07 lower accuracy than GLM-4-0414 on NonambigQA, yet it obtains a 0.037 lower ECE and a 0.078 higher AUROC. This indicates that accurate confidence estimation is not merely a byproduct of better task performance. Additionally, this highlights a potential motivation for using reasoning models even on simple factual question-answering tasks: although they may not achieve higher accuracy, they communicate their confidence more reliably.
Interestingly, Qwen3 exhibits a substantial difference in calibration depending on whether it engages in CoT through Thinking or Non-thinking mode. This suggests that the calibration observed in reasoning models primarily stems from their slow thinking CoT, rather than from inherent properties of the model itself. We further support this claim in Section 4.3, where we show that even non-reasoning models can improve calibration by engaging in slow thinking through in-context learning.
Figure 2 (left) illustrates the superior calibration of R1-Distill-Qwen over Qwen2.5-Instruct on TriviaQA. Surprisingly, for confidence levels above 60%, R1-Distill-Qwen exhibits near-perfect calibration, where its estimated confidence closely matches the actual accuracy (i.e., points lie near the line $y=x$, where $y$ is accuracy and $x$ is confidence). Plotting the sample frequency of the confidence bins (right of Figure 2) reveals that Qwen2.5-Instruct is highly overconfident, estimating either the 85% or 95% confidence bin for more than 80% of the total samples. In contrast, R1-Distill-Qwen produces estimations across a relatively diverse range of confidence levels, indicating less tendency to be overconfident. Nonetheless, there is room for improvement for R1-Distill-Qwen as well, as it rarely estimates confidence lower than 55%.
Reasoning models are also better calibrated on reasoning-intensive datasets (Table 2), showing that our findings generalize to more complex tasks. However, we observe a few cases where non-reasoning models outperform their reasoning counterparts in AUROC—specifically, GLM-4-0414 on both Math subsets, and EXAONE-3.5-Instruct on the Non-Math subset of SuperGPQA. Upon closer inspection, we find that the root cause lies in the multiple-choice format of MMLU-Pro and SuperGPQA, which implicitly provides models with cues about their correctness based on the available options. Non-reasoning models, exhibiting overconfidence, tend to predict 95% confidence when their answers appear among the multiple-choice options, but when they do not, they select the closest available option and assign a lower confidence of 85%. Relying on just two confidence bins, reinforced by implicit cues from the multiple-choice options, gives non-reasoning models an unearned advantage in discriminative power, which inflates their AUROC scores. Nonetheless, the non-reasoning models receive worse Brier Scores, as the metric, unlike AUROC, directly penalizes mismatches between predicted confidence and actual outcomes.
Table 2: Benchmark results on reasoning-intensive datasets. The full result with Accuracy and ECE is available in Appendix A.3.

### v2

@@ Abstract
Despite their strengths, large language models (LLMs) often fail to communicate their confidence accurately, making it difficult to assess when they might be wrong and limiting their reliability. In this work, we demonstrate that reasoning models that engage in extended chain-of-thought (CoT) reasoning exhibit superior performance not only in problem-solving but also in accurately expressing their confidence. Specifically, we benchmark six reasoning models across six datasets and find that they achieve strictly better confidence calibration than their non-reasoning counterparts in 33 out of the 36 settings. Our detailed analysis reveals that these gains in calibration stem from the slow thinking behaviors of reasoning models (e.g., exploring alternative approaches and backtracking) which enable them to adjust their confidence dynamically throughout their CoT, making it progressively more accurate. In particular, we find that reasoning models become increasingly better calibrated as their CoT unfolds, a trend not observed in non-reasoning models. Moreover, removing slow thinking behaviors from the CoT leads to a significant drop in calibration. Lastly, we show that non-reasoning models also demonstrate enhanced calibration when simply guided to slow think via in-context learning, fully isolating slow thinking as the source of the calibration gains. [^1] Our code is available at https://github.com/MattYoon/reasoning-models-confidence
@@
[…]
1 Introduction
A persistent weakness of large language models (LLMs) is their tendency to sound confident even when they are wrong (huang2025survey, ; xiong2024can, ). This overconfidence threatens their reliability, especially in high-stakes scenarios (pal2023med, ; zhang2023siren, ). Meanwhile, recent reasoning models like OpenAI’s o1 (jaech2024openai, ) and Deepseek-R1 (deepseekr1, ) have demonstrated strong problem-solving capabilities through chain-of-thought (CoT) (wei2022chain, ) eliciting slow thinking behaviors such as exploring alternative approaches and verifying their answers (min2024imitate, ; gandhi2025cognitive, ). Yet, it remains underexplored whether such slow thinking behaviors also help LLMs to “know what they know” (kadavath2022language, ) and accurately communicate the limits of their knowledge.
To this end, we present an extensive study of reasoning models’ ability to accurately estimate and express confidence in their output; a process known as verbalized confidence estimation (tian-etal-2023-just, ; geng-etal-2024-survey, ). We demonstrate that reasoning models exhibit superior confidence calibration compared to non-reasoning models, and that this improvement stems from slow thinking behaviors, which allow LLMs to dynamically adjust their confidence throughout the reasoning process. For instance, as shown in Figure 1, the model’s confidence increases when it first verifies its answer (“Oh, and he also created the council of 500”), and decreases when it considers an alternative (“Pericles came later”). The confidence then rises again as it rejects the alternative (“but I don’t think he was”), and reaches its highest level with further verification (“I don’t think I’m mixing up any details”).
We then conduct a detailed analysis to support our claim that the calibration gains in reasoning models stem from slow thinking, which allows them to dynamically adjust and correct their confidence throughout the reasoning process. (1) We measure changes in calibration metrics throughout the CoT process and observe a steady, gradual improvement for reasoning models, with statistically significant trends ($p<0.05$). Non-reasoning models show no significant trend and in some cases, surprisingly, even exhibit a worsening pattern. (2) In our ablation study, we find that the non-linear structure of slow thinking gandhi2025cognitive , such as the model’s ability to explore alternatives and revise its reasoning, plays a critical role in improving calibration. Conversely, prompting the model to explicitly reason about its own confidence yields only limited gains in calibration. (3) Lastly, we observe that non-reasoning models can also benefit from slow thinking when simply guided via in-context learning (dong2022survey, ), supporting our claim that these improvements arise from the slow thinking process itself, rather than from some other factors inherent to reasoning models.
@@ 2 Related Work
Various approaches have been proposed for confidence estimation, which is the process of inferring an LLM’s confidence in its predictions (geng-etal-2024-survey, ). This includes measuring token-level probabilities of answers (gupta2024language, ), training proxy probe models on internal hidden states (kadavath2022language, ), and applying supervised fine-tuning with labeled confidence data (xu2024sayself, ). Nonetheless, these methods assume access to a model’s internal states or weights, which are usually unavailable in proprietary LLMs (shrivastava2023llamas, ; xie2024survey, ). Additionally, they are often model-specific, necessitating readjustment or retraining when the underlying LLM changes (yang2024verbalized, ).
An alternative approach that avoids these limitations involves sampling multiple responses for the same prompt and deriving confidence based on response consistency (manakul-etal-2023-selfcheckgpt, ). While this method is compatible with proprietary LLMs and is model-agnostic, its primary limitation is the increased computational cost due to repeated inference (yang2024verbalized, ).
In this work, we focus on verbalized confidence estimation, where the LLM is prompted to directly express its confidence as a part of their output, either through a linguistic phrase (e.g., “Highly likely”) or a numerical probability (e.g., “0.95”) (geng-etal-2024-survey, ). Verbalized confidence estimation has emerged as a widely studied approach as it is compatible with proprietary LLMs, model-agnostic, and computationally efficient (tian-etal-2023-just, ; xiong2024can, ; yang2024verbalized, ). While verbalized confidence estimation offers ease of use and broad applicability, its accessibility also makes its limitations more consequential. Overconfidence is a persistent issue: LLMs tend to express high certainty even when they are incorrect, posing serious risks for real-world deployment of LLMs (xiong2024can, ; zhou2024relying, ). Our results suggest that reasoning models and their use of slow thinking hold strong potential to mitigate this issue.
@@
Increasing compute at inference time, or inference-time scaling, has become a focus for improving the reasoning performance of LLMs (snell2025scaling, ). One effective approach to inference-time scaling is training models to generate longer CoTs, as demonstrated by models like OpenAI’s o1 (jaech2024openai, ) and DeepSeek-R1 (deepseekr1, ), which are commonly referred to as reasoning models. The key factor that distinguishes the CoT of reasoning models from that of non-reasoning, instruction-tuned models beyond just length is their slow thinking (min2024imitate, ), which drives their problem-solving ability. Slow thinking mirrors human cognitive behaviors, characterized by non-linear reasoning traces such as exploring alternative approaches, verifying answers, and backtracking (gandhi2025cognitive, ). During slow thinking, reasoning models also frequently produce epistemic markers that signal uncertainty (e.g., “I think”, “maybe”) which are rarely observed in non-reasoning models (zhou2024relying, ). Structurally, reasoning models typically enclose their CoT within special tokens such as <think>...</think> to isolate the reasoning phase, in contrast to non-reasoning models.
@@ 3 Reasoning models better express their confidence
In this section, we demonstrate that reasoning models express their confidence more accurately than non-reasoning models, first detailing the experimental setup (Section 3.1) and then presenting the results (Section 3.2). Full results and additional experiments are provided in Appendix A, and further setup details are provided in Appendix B.
@@ 3.1 Experiment setup
We evaluate LLM calibration across two types of datasets: knowledge-focused and reasoning-intensive. We use the knowledge-focused datasets, TriviaQA and NonambigQA (joshi-etal-2017-triviaqa, ; min-etal-2020-ambigqa, ; kwiatkowski-etal-2019-natural, ), to provide a more controlled setting for comparing the calibration of reasoning and non-reasoning models, as both achieve similar accuracy given that CoT does little to help solve these tasks (sprague2024cot, ). Therefore, this setup ensures that any differences in calibration are not simply due to reasoning models being better at solving the task. In addition, since reasoning models are primarily intended for complex reasoning tasks, we also include reasoning-intensive datasets, MMLU-Pro and SuperGPQA (mmlupro, ; supergpqa, ), to ensure our findings generalize. These datasets are challenging multiple-choice benchmarks centered on knowledge-driven reasoning. We use two subsets of each reasoning dataset: a Math subset focused on arithmetic reasoning, and a Non-Math subset covering other types of reasoning. Due to our broad range of experiments, we uniformly sample 1,000 examples from each dataset or subset to keep the compute manageable. To assess the variability of this choice, we perform bootstrapping over multiple resampled subsets and report the standard deviation in Appendix A.2.4.
@@
We benchmark a diverse set of six reasoning models derived from four different 32B-scale LLMs against their non-reasoning counterparts. [^2] We discuss the results on different of model sizes in Section 5. Specifically, we evaluate: R1-Distill-Qwen, QwQ, and OR1-Preview reasoning models (deepseekr1, ; qwq32b, ; skywork-or1-2025, ) against Qwen2.5-Instruct (qwen25technicalreport, ); GLM-Z1-0414 against GLM-4-0414 (glm2024chatglm, ); EXAONE-Deep (exaonedeep, ) against EXAONE-3.5-Instruct (exaone35, ); and Qwen3 with Thinking Mode (abbreviated as Qwen3 Thinking) against Qwen3 with Non-thinking Mode (Qwen3 Non-thinking) (qwen3, ). Unlike the other pairs, which consist of separate model checkpoints, Qwen3 is a hybrid model that supports both reasoning-style CoT (Thinking Mode) and non-reasoning-style CoT (Non-Thinking Mode) via prompting. [^3] The Non-thinking mode is enabled by injecting an empty thinking block, <think></think>, at the beginning of the model’s response.
@@
In a single turn of conversation, we instruct the models to perform three steps using CoT: (1) Solution Reasoning, where the model reasons step-by-step to arrive at an answer to the given question; (2) Confidence Reasoning, where it evaluates its own confidence for that answer step-by-step; and (3) Confidence Verbalization, where it maps their confidence in one of ten bins, ranging from “Almost no chance (0–0.1)” to “Almost certain (0.9–1.0)”. Each bin includes both a linguistic descriptor (e.g., “Almost certain”) and its corresponding numerical probability (e.g., “0.9–1.0”), which are inspired by the approaches used in prior work (tian-etal-2023-just, ; xiong2024can, ). To ensure a fair comparison, we use the same instructions for both reasoning and non-reasoning models. The full prompt we use is included in Appendix B.2.
For reasoning models, we expect all three steps to be carried out within the thinking process <think>...</think>. However, we observe that some reasoning models (specifically R1-Distill, OR1-Preview, and GLM-Z1) rarely engage in Confidence Reasoning. [^4] To elaborate, their thinking process mostly only contains Solution Reasoning, although they consistently do output their Confidence Verbalization outside of the thinking termination token </think> without CoT. To address this, we force these models to include Confidence Reasoning within their thinking process by generating up to the </think> token, and replacing the token with “Okay, now let’s assess my overall thinking process so far step-by-step. I need to evaluate how likely my answer is correct.” and performing a second round of inference. We ablate the effect of this choice in Section 4.2, where we observe that it has negligible impact on calibration.
As part of the instruction, models are asked to format their final response as “Answer:ANSWER Confidence:CONFIDENCE” after completing all three steps. We apply rule-based filtering using regular expressions to extract the predicted answer and confidence. Answers are matched against the ground truth using GPT-4o mini (openai_gpt4omini, ), leveraging the prompt from OpenAI’s simple-evals codebase (openai_simpleevals, ). We describe the full procedure we use to ensure that all model responses contain extractable answers and confidence predictions in Appendix B.2.
@@
We use commonly adopted calibration metrics from prior work (geng-etal-2024-survey, ; tian-etal-2023-just, ; xiong2024can, ). Expected Calibration Error (ECE(naeini2015obtaining, ) measures the average discrepancy between accuracy and predicted confidence within each confidence bin, weighted by the number of samples in each bin. While ECE is an intuitive metric for measuring absolute calibration, it fails to capture calibration at the individual prediction level and does not account for the model’s discriminative ability, which refers to how well the model assigns higher confidence to correct predictions over incorrect ones (geng-etal-2024-survey, ). In contrast, AUROC (bradley1997use, ) captures the model’s discriminative ability by computing the probability that a randomly chosen correct prediction is assigned higher confidence than a randomly chosen incorrect one, but it does not measure absolute calibration. Brier Score (brier1950verification, ) quantifies the mean squared difference between predicted confidence and the true binary outcome, capturing both absolute calibration at the individual level and the discriminative ability.
Alternative prompting strategies and setups
To ensure the robustness of our findings, we experiment with a wide range of alternative prompting strategies and setups, and provide the results in Appendix A.2. (1) We explore different confidence expression styles such as providing only a linguistic descriptor without a probability, or outputting a numerical probability directly without binning. (2) Since some reasoning models perform two rounds of inference due to the forced Confidence Reasoning, we also evaluate a similar two-step sequential prompting setup for non-reasoning models. (3) We evaluate more advanced prompting strategies for non-reasoning models, as suggested by prior work, such as Top-K and Multi-step prompting (tian-etal-2023-just, ; xiong2024can, ). (4) While our main experiments use greedy decoding to avoid randomness and ensure reproducibility, we also test sampling-based decoding to confirm that our results generalize across decoding strategies. Across all setups, our findings remain consistent: reasoning models exhibit better calibration than their non-reasoning counterparts.
@@ 3.2 Experiment result
We observe that reasoning models consistently outperform their non-reasoning counterparts across all calibration metrics on knowledge-focused datasets (Table 1). Notably, reasoning models exhibit superior calibration even in cases where they underperform non-reasoning models in task accuracy. For example, GLM-Z1-0414 achieves 0.07 lower accuracy than GLM-4-0414 on NonambigQA, yet it obtains a 0.037 lower ECE and a 0.078 higher AUROC. This indicates that accurate confidence estimation is not merely a byproduct of better task performance. Additionally, this highlights a potential motivation for using reasoning models even on simple factual question-answering tasks: although they may not necessarily achieve higher accuracy, they communicate their confidence more reliably.
Interestingly, Qwen3 exhibits a substantial difference in calibration depending on whether it engages in CoT through Thinking or Non-thinking Mode. As Qwen3 is a single checkpoint, this suggests that the calibration observed in reasoning models primarily stems from their slow thinking CoT, rather than some other inherent differences between the models. We further support this claim in Section 4.3, where we show that even non-reasoning models can improve calibration by engaging in slow thinking through in-context learning.
@@
Figure 2 (left) illustrates the superior calibration of R1-Distill-Qwen over Qwen2.5-Instruct on TriviaQA. Surprisingly, for confidence levels above 55%,55555% corresponds to the “Better than even (0.5–0.6)” bin. Throughout the paper, we denote each bin by the average value of its numerical confidence range for brevity. R1-Distill-Qwen exhibits near-perfect calibration, where its estimated confidence closely matches the actual accuracy (i.e., points lie near the line $y=x$, where $y$ is accuracy and $x$ is confidence). Plotting the sample frequency of the confidence bins (right of Figure 2) reveals that Qwen2.5-Instruct is highly overconfident, estimating either the 85% or 95% confidence bin for more than 80% of the total samples. In contrast, R1-Distill-Qwen produces estimations across a relatively diverse range of confidence levels, indicating less tendency to be overconfident. Nonetheless, there is room for improvement for R1-Distill-Qwen as well, as it rarely estimates confidence lower than 55%.
Reasoning models are also better calibrated on reasoning-intensive datasets (Table 2), showing that our findings generalize to more complex tasks. However, we observe a few cases where non-reasoning models outperform their reasoning counterparts in AUROC; specifically, GLM-4-0414 on both Math subsets, and EXAONE-3.5-Instruct on the Non-Math subset of SuperGPQA. Upon closer inspection, we find that the root cause lies in the multiple-choice format of MMLU-Pro and SuperGPQA, which implicitly provides models with cues about their correctness based on the available options. Non-reasoning models, exhibiting overconfidence, tend to select the 95% confidence bin when their answers appear among the multiple-choice options, but when they do not, the models select the closest available option and assign a lower confidence of 85%. Relying on just two confidence bins, reinforced by implicit cues from the multiple-choice options, gives non-reasoning models an unearned advantage in discriminative power, which inflates their AUROC scores. Nonetheless, the non-reasoning models receive worse Brier Scores, as the metric, unlike AUROC, directly penalizes mismatches between predicted confidence and actual outcomes.
Table 2: Benchmark results on reasoning-intensive datasets. The full result with ECE and Accuracy is available in Appendix A.3.